### v2

# Trade-offs between structural richness and communication efficiency in music network representations

Abstract
Music is a structured and perceptually rich sequence of sounds in time with well-defined symbolic features, whose perception is shaped by the interplay of expectation and uncertainty. Network science offers a powerful framework for studying its structural organization and communication efficiency. However, it remains unclear how feature selection affects the properties of reconstructed networks and perceptual alignment. Here, we systematically compare eight encodings of musical sequences, ranging from single-feature descriptions to richer multi-feature combinations. We show that representational choices fundamentally shape network topology, the distribution of uncertainty, and the estimated communication efficiency under perceptual constraints. Single-feature representations compress sequences into dense transition structures that support efficient communication, yielding high entropy rates with low modeled perceptual error, but they discard structural richness. By contrast, multi-feature representations preserve descriptive detail and structural specificity, expanding the state space and producing sharper transition profiles and lower entropy rates, which leads to higher modeled perceptual error. Across representations, we found that uncertainty increasingly concentrates in nodes with higher diffusion-based centrality while their perceptual error remains low, unveiling an interplay between predictable structure and localized surprise. Together, these results show that feature choice directly shapes music network representation, describing trade-offs between descriptive richness and communication efficiency and suggesting structural conditions that may support efficient learning and prediction.
I Introduction
Music is a highly complex phenomenon built from low-level features that combine into patterns incorporating multiple structural aspects. Gaining insight into how these patterns emerge from basic musical events offers a useful perspective for understanding both the complexity of music and the perceptual processes involved in listening Koelsch (2014); Koelsch et al. (2019), while also providing a model system for studying how complex structure arises in sequential data more generally.
Musical sequences can be represented at different levels of detail by selecting subsets of features—such as pitch, interval, or duration—that preserve some distinctions while discarding others, an idea closely related to viewpoint-based approaches in music cognition Conklin and Witten (1995). To study how these representational choices shape the resulting musical structure, network representations can be employed, in which sequence elements are modeled as nodes and observed transitions between elements as edges. Prior work used pitch-based representations with octave and duration Liu et al. (2009); Ferretti (2017), pitch-class sets Gomez et al. (2014); serrà_corral_boguñá_haro_arcos_2012; Nardelli (2023); Manaris et al. (2005); Serra-Peralta et al. (2021), and chord-split representations Kulkarni et al. (2024); Di Marco et al. (2025), which resulted in different topologies and statistical insights del Río et al. (2008); Zanette (2006); Moss et al. (2019); Mehr et al. (2019); Rohrmeier and Cross (2008). Understanding how these representational choices alter network topology and information structure requires a systematic comparison across levels of detail, one that makes explicit the trade-off between complexity reduction and the retention of relevant structure Tishby et al. (2000).
A second key aspect of representational choice concerns not the amount of information detail it preserves, but how its organization alters the uncertainty in musical sequence continuation that humans perceive. In particular, recent work shows that communication efficiency under perceptual constraints–quantified by comparing the true uncertainty in a sequence with the uncertainty humans effectively exhibit–depends on how that uncertainty is organized Lynn et al. (2020a). Human responses to sequential input are limited by memory, noise, and representational cost Lynn et al. (2020b); Momennejad et al. (2017); Howard and Kahana (2002); Meyniel et al. (2016), so the way transitions are organized directly affects how easily the sequence can be processed. For instance, Lynn et al. Lynn et al. (2020b) showed, using reaction-time experiments with visual sequences, that even when two sequences have the same average uncertainty, differences in the organization of transitions can cause humans to react as if the uncertainty were higher, resulting in a perceptual error—a mismatch between the true transition statistics and the observer’s inferred model. That is, structural organization affects the amount of information effectively perceived.
Also, real networks across domains, such as word networks or citation networks, show consistently lower modeled perceptual error than their randomized counterparts, indicating that their transitions are organized in ways that improve communication efficiency Lynn et al. (2020b, a). Crucially, this efficiency seems to differ across representations within the same domain: networks built from content words exhibit lower modeled perceptual error than those built from grammatical words Lynn et al. (2020b), reflecting that representational choices shape transition regularities. Together, these findings suggest that each representation defines its own uncertainty landscape and the way a representation structures uncertainty has an influence on the information available to a listener, and thus implications for perceptual difficulty—precisely the question we address in this work.
Here, we systematically compare eight network representations of music that are drawn from and extend existing approaches in the literature. These vary in their level of detail, from single-feature encodings (pitch, interval, duration) to multi-feature combinations that incorporate octave or duration with pitch. We quantify how these representational choices reshape network topology and evaluate how efficiently each representation structures information by computing both the resulting entropy rate and the divergence from a perceptual-constraint estimate. Together, these measures locate each encoding along a trade-off between descriptiveness—the degree to which the underlying musical structure is preserved—and communication efficiency, the amount of information conveyed with minimal modeled perceptual error Tishby et al. (2000); Lynn et al. (2020b). In addition, by further examining how uncertainty and modeled error are distributed locally across each network representation—and how these patterns vary with piece length—we identify conditions under which random-walk centrality aligns with uncertainty. Such structures may facilitate downstream statistical compression or predictive abstraction Huron (2008); Pearce (2018), offering clues about why certain structural forms are easier for humans to learn. More broadly, this network-based framework links the information-theoretic costs of a representational choice to the structural patterns it reveals, offering a general approach for analyzing how complexity is organized across different representational levels and suggesting structural conditions that may support efficient learning and prediction.
II Results
Figure 1: Networks reconstructed from a musical piece. (a) A simple composition with highlighted interval changes. (b) Eight distinct network representations. All networks are directed and weighted: edges indicate transitions, and weights correspond to their frequencies.
To examine how different levels of musical description shape structural and informational properties, we constructed eight network representations of the same sequences, spanning both established encodings from the literature and controlled feature-level variants introduced for systematic comparison. We start with minimal, single-feature encodings that expose the basic building blocks of musical structure, and then introduce additional features stepwise, allowing us to observe how topological and informational patterns emerge or transform with increasing representational richness. In all cases, nodes correspond to unique feature configurations and edges record empirical transitions between consecutive elements in the musical sequence. The eight representations are defined as follows:
  • Pitch: Each node represents the note being played, identified by its pitch class (e.g., C, D, E), without distinguishing the octave. If several notes occur together (a chord), they form a single node serrà_corral_boguñá_haro_arcos_2012.
  • Duration: Each node represents how long a note or chord lasts, expressed as a standardized value relative to a beat.
  • Interval: Each node represents the distance in semitones between two consecutive notes or chords (see Fig. 1a).
  • Pitch+Duration: Each node represents the note or chord together with its duration.
  • Pitch+Octave: Each node represents the note or chord together with its octave information.
  • Pitch+Duration+Octave: Each node represents the note or chord with its duration and octave, combining all three features Ferretti (2017); Liu et al. (2009).
  • Pitch (split): A variation of the Pitch model in which chords are split, so that each note is represented as a separate node. Edges are created from the predecessor to each note in the chord, and from each note to the possible successors.
  • Pitch+Octave (split): A variation of the Pitch+Octave model in which chords are split, with each note and its octave treated as a separate node. As in the previous model, edges connect the predecessor to each note in the chord and from each note to the possible successors Kulkarni et al. (2024); Di Marco et al. (2025).
Given a musical piece in MIDI format, we constructed networks by tracking transitions between consecutive musical elements. Each transition forms a directed edge, while the definition of a node depends on the specific model chosen. The resulting networks are both directed and weighted, with edge weights corresponding to transition counts during the execution of the piece. Figure 1 illustrates the differences between the different representations. We detailed these differences in Supplementary Figure S1, which depicts network representations of Fryderyk Chopin’s Waltz in E-flat major, Op. 18.
In this work, we focus on piano compositions, which are both musically rich across multiple dimensions and well documented in large, curated corpora. Our corpus combines two sources: the piano-midi.de dataset Krueger (1996), which contains 268 pieces, and the MSDM dataset Dorfer et al. (2019), which contains 665 pieces. Together, these datasets cover works ranging from the Baroque period to the 20th century. We preprocessed and merged the two collections, separating the musical scores by left and right hands. Further details are provided in Methods Section IV.1.
II.1 Structural consequences of representational choice
Figure 2: Topological properties of music network models. Violin plots showing the distribution of eight network measures: (a) size NN, (b) average degree $\langle k\rangle$, (c) degree heterogeneity $\langle k^{2}\rangle/\langle k\rangle^{2}$, (d) average clustering coefficient $\langle C\rangle$, (e) average shortest path length $\langle l\rangle$, (f) reciprocity $r$, (g) average in-strength $\langle s_{in}\rangle$, and (h) betweenness centrality BCBC, across different models averaged over all musical pieces for the right hand. Models are grouped by the number of musical features used. Each symbol corresponds to an individual model and also highlights the median of each distribution.
To assess the impact of the different model choices on network topology, we computed a range of undirected, directed, and weighted network measures for each representation (Supplementary Tables S1–S2). Figure 2 summarizes eight representative metrics for the right-hand networks. The smallest networks arise in the Duration and Pitch(split) models (Fig. 2a), whereas the Pitch+Duration and Pitch+Octave+Duration abstractions produce the largest and sparsest structures, with correspondingly low average degree (Fig. 2b), clustering coefficient (Fig. 2d), reciprocity (Fig. 2f), and betweenness centrality (Fig. 2h). In contrast, the Pitch(split) and Pitch+Octave(split) models yield the highest degrees and reciprocity, as splitting chords inflates local branching and introduces many short cycles. These contrasts illustrate how more compressed representations tend to concentrate transitions into dense, uniform structures, whereas adding descriptive features expands the state space and reveals finer organizational detail. Degree heterogeneity (Fig. 2c) remains low for the split models but increases systematically as additional features are incorporated, reflecting a controlled expansion of the state space, where new nodes increase in a structured manner and organize into progressively more diverse structural roles. Average shortest-path length (Fig. 2e) scales with network size across representations, while average in-strength (Fig. 2g) decreases inversely, since the same number of transitions is redistributed across more nodes. Additional metrics for left-hand networks show the same patterns (Supplementary Figs. S2–S4).
Collectively, these findings show that representational choices do not simply scale the size of the network but reshape its structure in a systematic way. As additional musical features are incorporated, the state space expands in a way that networks become progressively more heterogeneous. This controlled expansion reflects how musical structure organizes itself across representational levels and highlights the distinctive regimes—dense, compressed configurations versus sparse and more heterogeneous ones—that different representations occupy.
II.2 Informational and perceptual effects of representational choice
Building on these topological patterns, we next examine how such structural differences influence the organization of musical information, shifting from network metrics to measures of information content. Following prior work Kulkarni et al. (2024), we examined how each representation organizes uncertainty by analyzing the Markov transition structure through the Shannon entropy of a random walk on the network Gómez-Gardenes and Latora (2008); Shannon (1948) (see Section IV.2). Entropy here serves as a global measure of how the network channels—and restricts—possible continuations, providing a complementary perspective to the topological results above. It is computed by weighting each node’s local entropy, i.e., the uncertainty in its outgoing transitions, by the node’s stationary distribution:
$S=\sum_{i}\pi_{i}S_{i}=-\sum_{i}\pi_{i}\sum_{j}P_{ij}\log P_{ij},$ (1)
where $\pi_{i}$ denotes the stationary visitation probability of node $i$. This probability is calculated as the entries in the normalized eigenvector of the transition matrix $\mathbf{P}$, formed by the probabilities $P_{ij}$, corresponding to the eigenvalue 1, see details in Methods Section IV.2. Because the stationary distribution reflects how often a diffusive random walker encounters each node in the long run, the resulting entropy captures how uncertainty is dynamically organized across the representation. In information-theoretic terms, it corresponds to the entropy rate or the average information produced per step of the encoded sequence.
Figure 3: Information content and perception efficiency in music network models. (a) Median entropy of weighted $S$ versus unweighted $S^{\mathrm{uw}}$ networks. (b) Median entropy $S$ versus entropy of type-A randomized networks. (c) Median Kullback–Leibler divergence of weighted $D_{KL}$ versus unweighted $D^{\mathrm{uw}}_{KL}$ networks. (d) Median $D_{KL}$ versus $D_{KL}$ of type-A randomized networks. (e) Median $D_{KL}$ versus median $S$. (f) Median $D^{\mathrm{uw}}_{KL}$ vs median $S^{\mathrm{uw}}$. Models are grouped by the number of musical features used. Each symbol corresponds to an individual model. Error bars show the interquartile range (IQR) around the median.
Figure 3a shows that entropy varies systematically across representations, with simpler models generally exhibiting higher uncertainty and richer representations yielding lower average information per step. Duration is a notable exception: despite its simplicity, it exhibits relatively low entropy because it has a very small vocabulary and low average degree, which strongly constrains the set of possible continuations. As additional features are introduced, the vocabulary expands, and the networks become sparser (see Fig. 2a,b, and Supplementary Figure S5). Although a larger state space could, in principle, increase uncertainty, the added specificity makes many transitions more predictable, thereby lowering the average information produced per step. Importantly, this entropy reduction does not reflect a homogeneous decrease in uncertainty across all states.
Richer representations develop a pronounced heterogeneity (Fig. 2c). From an information perspective, this implies that many nodes have near-deterministic successors, whereas others act as hubs with multiple plausible continuations. By contrast, chord-splitting models inject extra local branching, inflating uncertainty despite using the same underlying musical material. This effect is evident when comparing Pitch+Octave versus Pitch+Octave(split) models, where splitting chords yields substantially higher entropy. To distinguish uncertainty due solely to connectivity from uncertainty influenced by empirical transition frequencies (i.e., how often each transition occurs in the corpus), Figure 3a also compares entropy in weighted versus unweighted networks. In the weighted case, transition probabilities $P_{ij}$ are estimated from observed counts, whereas in the unweighted case we set all outgoing edges from a node to be equiprobable ($P_{ij}=1/k_{i}^{out}$), recompute the stationary distribution $\pi$, and then evaluate Eq. (1). The largest discrepancy appears in the Duration model, where ignoring weights makes this highly compressed representation appear far more uncertain than it actually is.
To separate uncertainty arising from local node-level constraints from that produced by the global arrangement of transitions, we compared each representation with a degree- and out-strength–preserving null model (Rand A), which fixes the local entropies at each node. Any differences in global entropy therefore reflect how the stationary distribution reallocates weight across these unchanged local uncertainties. Across representations—except for the most compressed ones (Duration and Pitch(split)), where structure is minimal—real networks consistently exhibit higher entropy than their randomized counterparts (Fig. 3b). This shows that visitation tends to concentrate in high-uncertainty regions, increasing the information produced per step beyond what the same local constraints would predict.
A second null model (Rand B), which randomizes edge weights while preserving the network skeleton, yields a comparable entropy reduction (Supplementary Figs. 6-7), confirming that the specific placement of transition frequencies—not only the topology—plays a crucial role in shaping how uncertainty is distributed across a representation.
Uncertainty is tightly linked to the amount of perceptual effort involved in correctly tracking a musical sequence. Next, we quantify the perceptual difficulty of each representation, enabling us to assess how representational choices affect access to the encoded musical information. We use the framework developed by Lynn et al. in Lynn et al. (2020b), which introduced a feature-agnostic model that estimates computationally the perceived transition structure under cognitive constraints. The model produces an inferred transition matrix $\hat{\mathbf{P}}$ from the true transition matrix,
$\mathbf{\hat{P}}=(1-\eta)\,\mathbf{P}\,\left(\mathbf{I}-\eta\,\mathbf{P}\right)^{-1},$ (2)
which deviates from the true probabilities in ways consistent with empirically observed limitations due to inaccuracies, controlled by parameter $\eta$, and the integration of higher-order paths. To measure the deviation, we computed the KL divergence between the inferred transition matrix $\mathbf{\hat{P}}$ and the true matrix $\mathbf{P}$ across representations. Lower divergence indicates that the perceptual model reconstructs the transition structure more faithfully, and thus that the representation communicates information more efficiently. See Methods Section IV.3 for more details.
KL divergence is lowest for single-feature models such as Pitch or Duration, and increases as additional features are combined, reaching its maximum in the Pitch+Octave+Duration model (Fig. 3c). In other words, simpler representations appear more perceptually efficient—they produce transition structures that are easier for the model to infer—even though they omit musically relevant distinctions. The strong performance of split models, however, should be interpreted with caution, as their artificially increased local branching makes them less sensitive to the smoothing applied in the perceptual estimate. Both null models confirm this pattern: for Rand A, KL divergence is consistently higher than in the real networks (Fig. 3d), except in the smallest representations, where real and randomized sequences are almost indistinguishable; Rand B yields a similar inflation (Supplementary Figures 6-7). Moreover, Figure 3c shows that, in line with results from Kulkarni et al. Kulkarni et al. (2024), incorporating edge weights consistently reduces KL divergence, confirming that empirical transition probabilities improve the alignment between the perceptual model and the true structure. This effect is most pronounced in the Duration representation, where weights seem to capture essential temporal information.
Taken together, the topological and information-theoretic results reveal a coherent progression across representational levels (Fig. 3e,f). Highly compressed, single-feature models produce smaller, denser networks with relatively uniform branching—a structure that yields high entropy per step but also low KL divergence. These representations act as efficient communication channels: their limited state space and dense topology allow the perceptual model to reconstruct transition patterns with minimal error, even though the musical distinctions they encode are coarse. An exception is the Duration model, whose extremely small state space combined with heterogeneous edge weights yields low entropy, and whose KL divergence remains low because smoothing alters only a very limited set of transitions. As additional features are incorporated, network size grows, and connectivity becomes sparser and more heterogeneous. Degree variation increases, locally deterministic states emerge, and hubs with multiple continuations appear—topologies that concentrate information unevenly across the network. This controlled expansion lowers entropy because transitions become more specific, yet it simultaneously raises KL divergence, as the perceptual model struggles to approximate sharply peaked transition profiles. Therefore, the very structural refinement that enriches the representation creates a more challenging information landscape.
In combination, these patterns show that musical representations occupy a continuum between efficiency and descriptive detail. Compressed networks communicate information effectively but flatten musically relevant structure, whereas multi-feature networks preserve that structure at the cost of greater perceptual error. The topology of each representation—its size, branching patterns, and heterogeneity—largely determines where it sits along this continuum, clarifying how the choice of features reshapes both the organization of musical information and its accessibility under perceptual constraints.
II.3 Alignment of uncertainty and inference across musical representations
Building on the comparison of network representations, we now investigate how uncertainty and perception efficiency are organized within these networks. Specifically, we ask how entropy and perceptual errors are distributed along the network, how they vary with piece length, and how they manifest at the local level.
Figure 4: Global alignment of uncertainty and inference in music networks. (a) Median entropy $S$ versus mean node-level entropy $\bar{S}$. (b) Median KL divergence $D_{KL}$ versus node-averaged KL divergence $\bar{D}_{KL}$. (c) Mean node-level entropy with and without edge weights ($\bar{S}$ versus $\bar{S}^{\mathrm{uw}}$). (d) Median node-averaged KL divergence with and without edge weights ($\bar{D}_{KL}$ versus $\bar{D}^{\mathrm{uw}}_{KL}$). (d) Median entropy $S$ versus the difference between median and node-level entropy ($S-\bar{S}$). (e) Median KL divergence $D_{KL}$ vs the difference between median and node-averaged KL divergence ($D_{KL}-\bar{D}_{KL}$). In (d) and (e) data is grouped by transition count. Transition counts are divided into five equal-sized bins. Smaller and more transparent markers correspond to shorter pieces. Points show medians, and error bars indicate the interquartile range. All panels correspond to right-hand tracks. Models are grouped by the number of musical features used. Each symbol corresponds to an individual model.
We first compare global entropy $S$ with the average node entropy $\bar{S}=\frac{1}{N}\sum_{i}S_{i}$, which reflects node-level uncertainties without weighting by the stationary distribution. Across all models, $\bar{S}$ is consistently lower than $S$ (Fig. 4a), showing that high-entropy nodes receive more stationary mass. The increase in global entropy arises not only from degree heterogeneity but also from the alignment between stationary random-walk visitation and local uncertainty. This is evident from the out-strength–preserving null model (Rand A), which maintains the same heterogeneity yet shows lower entropy because this alignment is reduced (Fig. 3b). As a result, informational complexity becomes concentrated in a small set of frequently visited nodes. Left-hand results follow the same pattern (Supplementary Figure 6-7).
Next, we examine whether the inference model shows a similar alignment. The node-averaged KL divergence $\bar{D}_{KL}$, which ignores the stationary distribution (see Eq. 6), is consistently higher than the global $D_{KL}$ (Fig. 4b). Weighted and unweighted networks display nearly identical $\bar{D}_{KL}$ values (Fig. 4d), showing that weights do not improve inference at individual states. Rather, they reshape the stationary distribution so that transitions easier for the inference model to approximate are visited more often. This pattern mirrors the entropy results and shows that stationary random-walk visitation naturally concentrates on transitions whose statistical structure aligns better with the perceptual model.
Then, we examine how this alignment varies with piece length. Global entropy increases with the number of transitions in a piece (Fig. 4e), whereas $\bar{S}$ remains largely stable (Supplementary Figure 8a). Consequently, the gap $S-\bar{S}$ increases with length (Fig. 4e), meaning that high-entropy nodes absorb more stationary mass as pieces become longer. Chord-splitting models (Pitch(split), Pitch+Octave(split) show weaker alignment and a completely different evolution since uniformly high branching flattens the entropy landscape, and the Duration model shows a similar pattern due to its reduced structure.
The KL divergence shows a parallel trend. Global $D_{KL}$ decreases slightly with length (Fig. 4f), while $\bar{D}_{KL}$ remains stable (Supplementary Figure 8b), so the gap $D_{KL}-\bar{D}_{KL}$ increases (Fig. 4f). This indicates that longer pieces direct flow toward transitions that the inference model approximates well. The Pitch(split) model again deviates: its extreme compression causes networks to approach full connectivity as length increases, reducing structural distinctions and breaking the alignment between inference error and flow. In these cases, uncertainty remains high but loses meaningful organization, illustrating again how some modeling choices can result in degenerated structures. The two previous null models confirm that randomized networks exhibit weaker alignment and reduced length-dependent variation than the real data (Supplementary Figures 9-10).
Finally, to provide a more detailed perspective, we examine how these patterns manifest at the local level. To do so, we grouped nodes into quartiles according to their cumulative stationary distribution ($\pi$), with low-quartile nodes visited infrequently and high-quartile nodes visited more often. As shown in Figs. 5a,b, both short and long pieces exhibit a skewed visitation pattern: most nodes are rarely visited, while a small subset concentrates the majority of visitation mass. This effect becomes more pronounced in longer pieces and in multi-feature representations. Average node entropy $\bar{S}$ is higher in high-$\pi$ quartiles (Figs. 5c,d), while node-averaged KL divergence $\bar{D}_{KL}$ is minimized in the same regions (Figs. 5e,f). This pattern confirms that uncertainty is concentrated in central, high-$\pi$ nodes, while inference errors are relegated to the periphery. Because the stationary distribution is skewed, low-error central nodes dominate the overall inference, and high-error nodes carry little weight, demonstrating that the network structure itself mitigates the impact of the errors. This effect becomes even stronger in longer pieces, where the skew is more pronounced.
Figure 5: Local alignment of uncertainty and inference in music networks. (a,b) Median fraction of nodes required to accumulate successive quartiles of stationary mass ($M$). (c,d) Average node entropy ($\bar{S}$) across $\pi$ quartiles.(e,f) Node-averaged KL divergence ($\bar{D}_{KL}$) across $\pi$ quartiles. Points denote medians and error bars show interquartile ranges. Columns separate short and long musical compositions, defined by dividing pieces into two equal bins by the number of transitions. All panels correspond to right-hand tracks. Models are grouped by the number of musical features used. Each symbol corresponds to an individual model.
Overall, these results show that musical networks align uncertainty and effective inference with the visitation frequency, and that this alignment strengthens with piece length. As additional features expand the state space, these landscapes become more pronounced: uncertainty concentrates in dynamically central states while inference errors diminish in regions carrying more weight. This progression reflects the emergence of efficient, heterogeneous structures that maintain coherence as representational detail increases and, in doing so, provide a landscape that facilitates the extraction of recurrent patterns.
III Discussion
Representational choice plays a decisive role in the structural properties of musical networks. Our results show that the number and grouping of features—and whether musical events are split or not—substantially influence the resulting topology. Properties often associated with musical corpora, such as small-world organization Ferretti (2017); Liu et al. (2009), heavy-tailed degree distributions, or Zipf- and Heaps-like scaling Manaris et al. (2005); Zanette (2006), emerge only under certain representational levels and are not inherent to the raw data alone. More detailed, multi-feature representations tend to produce markedly heterogeneous and skewed structures, whereas compressed vocabularies attenuate these patterns. These findings show that topological regularities in music depend on how the underlying symbolic space is defined.
Building on this, the representational differences instantiate a broader trade-off between communication efficiency and descriptive richness. The limited perceptual error associated with compressed representations despite their high entropy implicates that information from higher-order transitions several steps apart in the musical sequence do not disturb substantially efficient communication, whereas higher-order transitions in richer representations capturing finer musical structure deviate more importantly from local transitions, which increases perceptual error and lowers entropy rate. This echoes principles from information bottleneck theory Tishby et al. (2000), minimum-description-length formulations Rissanen (1983), and probabilistic models of expectation Pearce and Wiggins (2012), where representations balance fidelity with cognitive cost. Within the network formalism, these trade-offs become explicit: the vocabulary used to encode a sequence not only determines the information carried per step but also its structural cost, captured by the KL divergence between the true transition matrix and the cognitively constrained approximation. Higher KL values indicate that expanded vocabularies introduce statistical structure that is harder for a human-like perceptual system to track, reflecting the additional cost imposed by finer-grained distinctions. By making the structure of information explicit—locally through uncertainty and globally through the stationary distribution and KL divergence–this framework offers a new perspective on how representational choices and perceptual limitations jointly constrain sequential processing.
Although our analysis remains descriptive rather than cognitive, the emergence of efficient communication channels in simpler representations parallels several known properties of auditory and musical processing. Neural pathways handle pitch, duration, and timbre through partially distinct streams Peretz and Coltheart (2003), and hierarchical integration Peretz and Zatorre (2005); Koelsch (2011) resembles the progressive increase in structural complexity we observe as more features are added. At the same time, human expectations depend on learning and exposure Pearce (2018); Daikoku et al. (2021), and thus cannot be reduced to the statistical transitions in a corpus. Our results, therefore, do not model cognition directly, but they provide a structural account of how musical information is organized at different levels of descriptiveness—an organization that perceptual systems must interact with, regardless of the specific statistical learning models involved.
This structural organization also offers insight into why some musical systems support pattern discovery more effectively than others. In richer representations, uncertainty concentrates in dynamically central nodes while less central nodes remain relatively predictable. Such landscapes provide learners with stable contexts and informative deviations, a contrast to uniformly uncertain or near-random systems in which motifs and regularities become difficult to extract. Notably, this interplay between predictable structure and localized surprise has also been linked to affective responses in music, where heterogeneous uncertainty can enhance engagement and pleasure Huron (2008); Cheung et al. (2019); Gold et al. (2019). Although learning is beyond the scope of this work, the results suggest that differentiated uncertainty may facilitate both pattern extraction and the perception of structure in musical sequences.
Our representational hierarchy was chosen to build on established symbolic viewpoints in music-network research while enabling a controlled progression from coarse to detailed descriptions. Other musically informed viewpoints—such as scale-relative functions or learned feature spaces—may yield different structural signatures and should be explored in future work. Similarly, our focus on classical piano music provided a rich context for examining how complexity grows with representational detail, but the same framework can be applied to other repertoires. Our intuition is that genres with more repetitive or minimalist structures will show weaker differentiation across representational levels, offering a useful contrast to the dynamics observed here.
A natural extension of this work is to examine how expectations derived from statistical-learning models interact with the informational landscapes uncovered here. Models such as IDyOM Pearce and Wiggins (2012) introduce learned priors that may bias the inferred transition structure away from the raw distribution of events, effectively modifying perceived topology and uncertainty. Understanding how information at different representational levels contributes to these expectations—and how listeners integrate or prioritize these layers—could provide deeper insight into the efficient processing of musical complexity.
More broadly, this approach demonstrates how the structural and informational consequences of representational choice can be analyzed in any sequential domain. Music serves as a clear testbed because its feature combinations and hierarchical abstractions are well characterized, but the framework is easily transferable to other sequences and systems. By tracing how topologies and informational landscapes evolve with representational detail, we gain a principled way to compare how different feature sets shape the organization of sequential information and the emergence of efficient heterogeneous structures.
IV Methods
IV.1 Datasets and Preprocessing
We used two publicly available collections of piano music in MIDI format: the piano-midi.de dataset Krueger (1996) (268 pieces) and the MSDM dataset Dorfer et al. (2019) (665 pieces). After merging, the corpus comprised 933 files spanning works from the Baroque to the 20th century. Some duplicates remain, since different encodings of the same composition may capture distinct transcription choices.
MIDI files were selected only if they included tempo and time-signature metadata to ensure accurate quantization of durations. We further restricted the corpus to files containing exactly two tracks, corresponding to the left and right hands of a piano score. Each track was converted into a directed and weighted network, resulting in two networks for each composition. To ensure a well-defined entropy with a unique stationary distribution, we restrict our analysis to the largest strongly connected component (LSCC) of each network. Moreover, to ensure that the resulting networks faithfully represented the musical pieces, we retained only LSCCs that contained at least 80% of the nodes, yielding more than 600 networks per model.
During preprocessing, each musical element was encoded according to the chosen network representation (see Section II). Symbolic features include pitch (frequency, grouped into pitch classes an octave apart), duration (quantized relative to a quarter note = $1.0$), and chromatic interval (pitch distance in semitones between consecutive events). If multiple pitches occurred simultaneously, the highest pitch was used to compute intervals. Descriptive statistics on pitch ranges, note durations, chord distributions, and interval patterns for both datasets, including systematic differences between left and right hands, are provided in Supplementary Section IV.
IV.2 Entropy in complex networks
The Shannon entropy for a node $i$ can be written as Gómez-Gardenes and Latora (2008); Kulkarni et al. (2024)
$S_{i}=-\sum\limits_{j}P_{ij}\log P_{ij},$ (3)
where $P_{ij}$ are the entries of the transition matrix $\mathbf{P}$. In the case of directed and weighted networks, $P_{ij}=\omega_{ij}/s^{\mathrm{out}}_{i}$, where $\omega_{ij}$ is the edge weight and $s^{\mathrm{out}}_{i}$ is the out-strength of the node $i$. Here we consider only outgoing links from node $i$. In unweighted networks, the out-strength is replaced by the out-degree $k^{\mathrm{out}}_{i}$.
The entropy of the entire network has the form of Eq. 1. The stationary distribution $\pi$ satisfies the condition $\mathbf{P}\pi=\pi$ and can be obtained as the normalized eigenvector of $\mathbf{P}$ corresponding to the eigenvalue 1. For the simplest case, i.e., an undirected and unweighted graph, the stationary distribution becomes $\pi_{i}=k_{i}/2E$ where $k_{i}$ is the degree of node $i$ and $E$ is the total number of edges. It is worth noting that the formula is only applicable when a network has a single strongly connected component. In the preprocessing step, we ensure that we work with such networks.
IV.3 Human perception model
We adopt the human perception model introduced by Lynn et al. Lynn et al. (2020a), which we briefly summarize here for completeness, as it serves as the basis for our perceptual analysis. In this model, the inferred transition probabilities $\mathbf{\hat{P}}$ are expressed in terms of the true transition probabilities $\mathbf{P}$ as
$\mathbf{\hat{P}}=(1-\eta)\,\mathbf{P}\,\left(\mathbf{I}-\eta\,\mathbf{P}\right)^{-1},$ (4)
where $\eta\in[0,1]$ controls the inaccuracy of perceptual expectations, and $(\mathbf{I}-\eta\mathbf{P})^{-1}$ denotes a matrix inverse. This inverse can be expanded as a geometric series,
$(\mathbf{I}-\eta\mathbf{P})^{-1}=\sum_{t=0}^{\infty}(\eta\mathbf{P})^{t},$ (5)
which reveals that perception integrates not only direct (one-step) transitions but also higher-order paths—two, three, or more steps apart—each weighted by a decaying factor $\eta^{t}$. Consequently, $\mathbf{\hat{P}}$ represents a temporally smoothed higher-order estimate of the transition structure. Empirical evidence suggests an average human accuracy parameter of $\eta\approx 0.8$ which we use for our computation Lynn et al. (2020b).
To quantify the discrepancy in the perception of music networks, i.e., the difference between these two probabilities, we use the Kullback-Leibler (KL) divergence defined as:
$D_{KL}(\mathbf{P}\parallel\mathbf{\hat{P}})=-\sum_{i}\pi_{i}\sum_{j}P_{ij}\log\frac{\hat{P}_{ij}}{P_{ij}}.$ (6)
The lower the KL divergence, the smaller the difference between $\mathbf{P}$ and $\mathbf{\hat{P}}$. In addition, we define the uniform average of the KL divergence as
$\bar{D}_{KL}(\mathbf{P}\parallel\mathbf{\hat{P}})=-\frac{1}{N}\sum_{i=1}^{N}\sum_{j}P_{ij}\log\frac{\hat{P}_{ij}}{P_{ij}}.$ (7)
Acknowledgments
We thank Timoteo Carletti and Nicolás Bono Rosselló for helpful comments and feedback on the manuscript. L. B. R and R. J acknowledge support from the Bridge grant funded by Young Researchers of the Complex Systems Society (yrCSS). R. J. acknowledges support from the fellowship FI-SDUR funded by Generalitat de Catalunya. L. B. R acknowledges support from the ARIAC project (No. 2010235), funded by the Service Public de Wallonie (SPW Recherche). M. A. S and M. B. acknowledge support from grant PID2022-137505NB-C22 funded by MCIN/AEI/10.13039/501100011033 and by “ERDF/EU”. M. B. acknowledges the ICREA Academia award, funded by the Generalitat de Catalunya.
V Data availability
The network datasets used in this study are available from the sources referenced in the manuscript and the supplementary materials.
VI Code availability
To generate the different network models, we have produced software that allows us to read MIDI files from the given datasets, select the features that will be used to build the given model, and produce the graphs for each track in the MIDI file, together with a text file that corresponds to the given viewpoint sequence. We will provide an open-source code upon publication.